A rectangle has a perimeter of 284 kilometres and a height of 79 kilometres. What is the length of the base?
Answer by stanbon(75887)

A rectangle has a perimeter of 284 kilometres and a height of 79 kilometres. What is the length of the base?
------------
P = 2(L + W)
284 = 2(L + 79)
------
142 = L + 79
L = 63 km
--------------------
Cheers,
Stan H.
------------